# On the Ricci flow on Trees

### v2

@@ 1. Introduction
$\frac{\partial{w}_{e}(t)}{\partial t}=-\kappa_{e}(t){w}_{e}(t),$ (1)
@@
$\frac{\partial w_{e}(t)}{\partial t}=-\kappa_{e}(t)w_{e}(t)+w_{e}\sum\limits_{h\in E(G)}\kappa_{h}w_{h}(t),$ (2)
@@ 2. Preliminaries and Definitions
$w_{e}(t)=\frac{\tilde{w}_{e}(t)}{\sum_{h\in E}\tilde{w}_{h}(t)}.$ (3)
@@
$g(x)=\begin{cases}w_{ux}(t)+w_{uv}(t)&\text{if}~~x\sim u,x\neq v;\\ w_{uv}(t)&\text{if}~~x=u\\ 0&\text{if}~~x=v\\ -w_{vy}(t)&\text{if}~~y\sim v,y\neq u;\\ 0&\text{otherwise}.\end{cases}$
@@
$\kappa_{u\to v}=\frac{2w_{uv}\gamma(w_{uv})-\sum\limits_{x\sim u}w_{ux}\gamma(w_{ux})}{w_{uv}D_{u}}.$ (4)
@@
    This is a discrete analog of the Gauss-Bonnet theorem on surfaces on a finite tree. These suggest that the Ricci flow system could potentially achieve equilibrium as $t$ goes to infinity.
@@
    $\begin{split}\sum_{h\in E(T)}\kappa_{h}w_{h}&=\sum_{i\in V(T)}\frac{(2-d_{u})d_{u}}{D_{u}}\\ &=\sum_{\text{leaf~}e}w_{e}+\sum_{u\in V,d_{u}\geq 2}\frac{(2-d_{u})d_{u}}{D_{u}}.\end{split}$ (5)
@@
    $\prod\limits_{uv\in E(T)}w_{uv}(t)=\Big(\prod\limits_{uv\in E(T)}w_{uv}(0)\Big)\times{\rm e}^{-2t}.$ (6)
@@
$w_{uv}(0)+\frac{1}{2}w_{ux}(0)-\frac{1}{2}w_{ux}(0)e^{-\frac{2}{3}t}\leq w_{uv}(t)\leq w_{uv}(0)+w_{ux}(0)+w_{ux}(0)e^{-\frac{1}{2}t}.$
@@ 2.4. Some lemmas
Let $f$ be a $C^{1}$ function on $[a,b]$. If $f(a)<0$, and when $f=0$ we have $df/dt\leq 0$, then $f(b)\leq 0$.
@@ 3. Asymptotic behavior of weights under the Ricci flow
Since $f$ is not a leaf, $d_{v}\geq 2$. The term $w_{e}+\frac{d_{v}-2}{D_{v}}$ remains strictly positive throughout finite time periods, and can only approach zero at $t=\infty$. Thus, $w_{f}(t)-w_{e}(t)$ is an increasing monotonic function, then $\lim\limits_{t\to\infty}(w_{f}(t)-w_{e}(t))$ exists or is infinity. If $\lim\limits_{t\to\infty}(w_{f}(t)-w_{e}(t))=\infty$, clearly, the result of (ii) follows. If $\lim\limits_{t\to\infty}(w_{f}(t)-w_{e}(t))=\infty$ is finite, then
@@
(iii) $\frac{w_{e}(t)}{w_{g}(t)}$ converges monotonically to $1$ as $t\to\infty$.
@@
Let the metric $w$ of the tree $T$ evolve under the unnormalized Ricci flow 1. Let $e$ represent a leaf edge incident to vertex $u$ where $d_{u}\geq 3$. If $e$ attains the maximum initial value $w_{e}(0)$, then $\kappa_{e}(t)\geq\kappa_{g}(t)$ for any other leaf $g$ incident to $u$, and $\kappa_{e}(t)>1-\frac{d_{u}-2}{d_{u}^{\prime}}$ for $t\in[0,\infty)$.
@@
Firstly, consider the case when $\lim\limits_{t\to\infty}w_{f}$ is finite for some internal edge $f$, which implies, according to part (ib) of Proposition 1, that $\lim\limits_{t\to\infty}\kappa_{f}=0$. By the expression for $\kappa_{f}$, there exists an edge $h$ such that the ratio $\frac{w_{f}}{w_{h}}$ tends to infinity. Based on previous results, it follows that leaf $e$ is such an edge, then $w_{e}$ approaches zero. The results of (i) and (ii) follow.
@@
    $\kappa_{e}(t)$ $>1-\frac{d_{u}-2}{d_{u}-2}>0,$
    which indicates $\kappa_{e}$ decreases to zero, then the derivative $w^{\prime}_{e}=-\kappa_{e}w_{e}$ is negative and eventually tends to zero. Therefore, $w_{e}$ is monotonically decreasing and, being limited below by zero, it converges to a bounded value.
@@
Let the metric $w$ of T evolve under the Ricci flow (1). For a leaf $e$, if $w_{e}(t)$ tends to infinity, then there exists a time $t_{0}$ such that $\kappa_{e}(t)<0$ for all $t\geq t_{0}$. That is, $w_{e}(t)$ is strictly increasing for $t\geq t_{0}$.
@@
$\kappa_{e}^{\prime}(t)=(1-\frac{d_{u}-2}{w_{e}D_{u}})^{\prime}=\frac{d_{u}-2}{w_{e}^{2}D_{u}^{2}}(D_{u}^{\prime}w_{e}+D_{u}w_{e}^{\prime})=\frac{d_{u}-2}{w_{e}D_{u}}(\frac{\sum\limits_{v\sim u}\frac{\kappa_{uv}}{w_{uv}}}{\sum\limits_{v\sim u}\frac{1}{w_{uv}}}-\kappa_{e}).$ (7)
@@
Given that $w_{e}(t)$ increases without bound, there exists a sufficiently large time $t_{0}>0$ where $w_{e}^{\prime}(t_{0})>0$. Assume $w_{e}(t)$ does not have a monotonic increase on the interval $t_{0},\infty)$, then there exists $t_{1}>t_{0}$ such that $w_{e}^{\prime}(t_{1})<0$. By the continuity, there exists a point $a$ in $(t_{0},t_{1})$ such that $w_{e}^{\prime}(a)=0$. Given that $w_{e}(t)$ is always positive, we have that $\kappa_{e}(t_{0})<0$ and $\kappa_{e}(a)=0$.
@@
Notice that for any internal edge $uv$ and finite time $a$, we have $\kappa_{uv}(a)<0$. If $d_{u}^{\prime}=1$, there is only one leaf $e$, apparently $\kappa_{e}^{\prime}(a)<0$. Let $d_{u}^{\prime}>1$. W.l.o.g., we assume that $e$ attains the maximum initial weight among all leaves sharing $u$. Then for any leaf edge $g\sim e$, $\kappa_{g}(a)\leq 0$ according to Lemma 5. Since $D_{u}(t)$ is always positive, we conclude that
$\kappa_{e}^{\prime}(a)<0.$
@@
Similarly, for another leaf $g\sim e$, the condition $\kappa_{g}(t)\leq\kappa_{e}(t)$ ensures that $\kappa_{g}(t)\leq 0$ for $t>t_{0}$.
Hence, for every leaf $e$, if its unnormalized weight grows to infinity, there exists some time $t_{0}$ such that $w_{e}^{\prime}(t)\geq 0$ for all $t>t_{0}$. Next, we prove that $w_{e}^{\prime}(t)>0$. Note that $\kappa_{e}(t)$ cannot be zero at some point $t>t_{0}$, as it would imply $\kappa^{\prime}_{e}(t)<0$, which is impossible. The proof is complete.
@@ 4. Asymptotic behavior of curvature
Let the metric $w$ of T evolve under the Ricci flow (1). Consider $e$ as a leaf edge connected to vertex $u$ where $d_{u}\geq 3$, and $w_{e}(0)\geq w_{g}(0)$ for any leaf $g$ incident to $u$. Consequently, the following statements are valid.
(i) when $d_{u}^{\prime}=d_{u}-1$, $\lim\limits_{t\to\infty}\kappa_{e}(t)=\frac{1}{d_{u}-1}$;
@@
When $d_{u}^{\prime}=d_{u}-1$, according to Lemma 5, there exists $t_{0}>0$ such that, $\kappa_{e}(t)\geq 1-\frac{d_{u}-2}{d_{u}-1}=\frac{1}{d_{u}-1}$ for $t\geq t_{0}$, which gives
@@
We proved the result (i).
@@
On the other hand, suppose that $d_{u}^{\prime}=d_{u}-2$. Then, for sufficiently large $t$, we have $\kappa_{e}(t)\geq 0$, which implies that $w_{e}(t)$ is non-increasing. Consequently, for every internal edge $f\sim e$, the ratio $\frac{w_{e}}{w_{f}}$ is decreasing since the denominator is non-decreasing. It follows that
@@
which implies that $\frac{2-d_{u}}{w_{f}D_{u}}$ does not vanish along any time sequence, and hence neither does $\kappa_{f}$. On the other hand, since $\kappa_{e}$ remains non-negative for sufficiently large $t$, this contradicts the fact that $(\kappa_{f}-\kappa_{e})$ tends to zero.
Therefore, $\lim_{t\to\infty}\frac{w_{e}}{w_{f}}=0$ for every internal edge, it follows that
@@
Let the metric $w$ of tree T evolve under the Ricci flow (1). Let $e=uv$ be a leaf with $d_{u}\geq 2$. Then the unnormalized weight $w_{e}(t)$ is bounded if and only if $\kappa_{e}(t)$ tends to $\frac{1}{d_{u}-1}$ or zero.
Further, if $e$ attains the largest initial weight among all leaves sharing vertex $u$, the unnormalized weight $w_{e}(t)\downarrow 0$, that is, decreases after sufficiently large time.
@@
By Proposition 3, the limit $\lim_{t\to\infty}w_{e}(t)$ exists. Suppose that $\lim_{t\to\infty}w_{e}(t)<\infty$. By Barbalat’s Lemma, the derivative
@@
must then converge to zero. However, since $\kappa_{e}(t)$ does not approach zero along at least one time sequence (because the condition $d^{\prime}_{u}\leq d_{u}-2$ does not satisfy Proposition 5), it follows that $w_{e}(t)\to 0$ along this sequence. Since the limit of $w_{e}(t)$ exists, we conclude that in fact
@@
For each internal edge $f\sim e$, the ratio $\frac{w_{e}}{w_{f}}$ tends to zero as $t\to\infty$. Assume $d^{\prime}_{u}\leq d_{u}-3$, then we have
@@
Thus, if $\lim\limits_{t\to\infty}w_{e}(t)$ remains bounded, then $d^{\prime}_{u}$ is at least $d_{u}-2$ and $\kappa_{e}(t)$ approaches either $\frac{1}{d_{u}-1}$ or zero.
@@
Thus, the derivative $w^{\prime}_{e}(t)$ is negative for large time, the unnormalized weight $w_{e}(t)$ decreases after sufficiently large time. Then $\lim\limits_{t\to\infty}w_{e}(t)$ is finite.
@@
$\kappa_{e}\leq 1-\frac{d_{u}-2}{d_{u}^{\prime}+\frac{(d^{\prime\prime}_{u}-2)d^{\prime\prime}_{u}}{(d^{\prime\prime}_{u}-2+d_{u})}}.$ (8)
@@
According to Proposition 5, if $d^{\prime}_{u}\geq d_{u}-2$, as well as $d^{\prime}_{v}\geq d_{v}-2$, then there exist leaves attached to $u$ and $v$ such that the ratio of their unnormalized weights to the unnormalized weight of the edge $f$ approaches zero. Consequently, $\kappa_{f}$ tends to zero as $t$ approaches infinity as both $\kappa_{u\to f}\to 0$ and $\kappa_{v\to f}\to 0$ hold. ∎
@@ 5. The limit metric with constant curvature zero
Proof of result (3) Let $L$ the set of leaf edges incident to $v_{1}$ or $v_{\ell}$. For finite $t$,
$\sum\limits_{\text{leaf~}e\sim v_{1}}\kappa_{e}(t)$ $>d_{v_{1}}-1-\frac{d_{v_{1}}-2}{D_{v_{1}}}\sum\limits_{\text{leaf~}e\sim v_{1}}\frac{1}{\tilde{w}_{e}}$
@@
$\sum\limits_{e\in L}\kappa_{e}(t)\downarrow 2,\quad t\to\infty.$
@@
The initial metric influences the limiting weight value for a caterpillar tree. Consider a caterpillar tree where each spine vertex has a degree of $2$. Thus, leaf edges only incident to the endpoints of the main path.
@@
For trees that are not of the caterpillar type, the unnormalized weight can approach infinity in the limit, a result of the number of leaves having curvature values that approach positive values.
@@
Repeating this argument for at least three different branches from $v_{0}$ shows that there are at least three such internal vertices in $T$. ∎

### v3

@@ 1. Introduction
(1) $\frac{\partial{w}_{e}(t)}{\partial t}=-\kappa_{e}(t){w}_{e}(t),$
@@
(2) $\frac{\partial w_{e}(t)}{\partial t}=-\kappa_{e}(t)w_{e}(t)+w_{e}\sum\limits_{h\in E(G)}\kappa_{h}w_{h}(t),$
@@ 2. Preliminaries and Definitions
(3) $w_{e}(t)=\frac{\tilde{w}_{e}(t)}{\sum_{h\in E}\tilde{w}_{h}(t)}.$
@@
$g(x)=\begin{cases}w_{ux}+w_{uv}&\text{if}~~x\sim u,x\neq v;\\ w_{uv}&\text{if}~~x=u\\ 0&\text{if}~~x=v\\ -w_{vy}&\text{if}~~y\sim v,y\neq u;\\ 0&\text{otherwise}.\end{cases}$
@@
(4) $\kappa_{u\to v}=\frac{2w_{uv}\gamma(w_{uv})-\sum\limits_{x\sim u}w_{ux}\gamma(w_{ux})}{w_{uv}D_{u}}.$
@@
    This is a discrete analog of the Gauss-Bonnet theorem on a finite tree. These suggest that the Ricci flow system could potentially achieve equilibrium as $t$ goes to infinity.
@@
    (5) $\begin{split}\sum_{h\in E(T)}\kappa_{h}w_{h}&=\sum_{u\in V(T)}\frac{(2-d_{u})d_{u}}{D_{u}}\\ &=\sum_{\text{leaf~}e}w_{e}+\sum_{u\in V,d_{u}\geq 2}\frac{(2-d_{u})d_{u}}{D_{u}}.\end{split}$
@@
    (6) $\prod\limits_{uv\in E(T)}w_{uv}(t)=\Big(\prod\limits_{uv\in E(T)}w_{uv}(0)\Big)\times{\rm e}^{-2t}.$
@@
$0<w_{uv}(t)\leq w_{uv}(0)+w_{ux}(0)-\frac{1}{2}w_{ux}(0)e^{-\frac{1}{2}t}.$
@@ 2.4. Some lemmas
Let $f$ be a $C^{1}$ function on $[a,b]$. If $f(a)<0$, and when $f=0$ we have $df/dt<0$, then $f(b)\leq 0$.
@@ 3. Asymptotic behavior of weights under the Ricci flow
Since $f$ is not a leaf, $d_{v}\geq 2$. The term $w_{e}+\frac{d_{v}-2}{D_{v}}$ remains strictly positive throughout finite time periods, and can only approach zero at $t=\infty$. Thus, $w_{f}(t)-w_{e}(t)$ is an increasing monotonic function, then $\lim\limits_{t\to\infty}(w_{f}(t)-w_{e}(t))$ exists or is infinity. If $\lim\limits_{t\to\infty}(w_{f}(t)-w_{e}(t))=\infty$, clearly, the result of (ii) follows. If $\lim\limits_{t\to\infty}(w_{f}(t)-w_{e}(t))$ is finite, then
@@
(iii) $\frac{w_{e}(t)}{w_{g}(t)}$ converges to $1$ as $t\to\infty$.
@@
Let the metric $w$ of the tree $T$ evolve under the unnormalized Ricci flow (1). Let $e$ represent a leaf edge incident to vertex $u$ where $d_{u}\geq 3$. If $e$ attains the maximum initial value $w_{e}(0)$, then $\kappa_{e}(t)\geq\kappa_{g}(t)$ for any other leaf $g$ incident to $u$, and $\kappa_{e}(t)>1-\frac{d_{u}-2}{d_{u}^{\prime}}$ for $t\in[0,\infty)$.
@@
Firstly, consider the case when $\lim\limits_{t\to\infty}w_{f}$ is finite for some internal edge $f$, which implies, according to part (ib) of Proposition 1, that $\lim\limits_{t\to\infty}\kappa_{f}=0$. Based on previous results, it follows that leaf $e$ is an edge such that $\frac{w_{f}}{w_{e}}\to 0$, then $w_{e}$ approaches zero. The results of (i) and (ii) follow.
@@
    $\kappa_{e}(t)$ $>1-\frac{d_{u}-2}{d_{u}-2}>0.$
    Thus, the derivative $w^{\prime}_{e}=-\kappa_{e}w_{e}$ is negative and then $w_{e}(t)$ is monotonically decreasing and, being limited below by zero, it converges to a bounded value.
@@
Let the metric $w$ of T evolve under the Ricci flow (1). For a leaf $e$, if $w_{e}(t)$ tends to infinity, then there exists a time $t_{0}$ such that $\kappa_{e}(t)<0$ for all $t\geq t_{0}$. That is, $w_{e}(t)$ is monotonically increasing for $t\geq t_{0}$.
@@
(7) $\kappa_{e}^{\prime}(t)=(1-\frac{d_{u}-2}{w_{e}D_{u}})^{\prime}=\frac{d_{u}-2}{w_{e}^{2}D_{u}^{2}}(D_{u}^{\prime}w_{e}+D_{u}w_{e}^{\prime})=\frac{d_{u}-2}{w_{e}D_{u}}(\frac{\sum\limits_{v\sim u}\frac{\kappa_{uv}}{w_{uv}}}{\sum\limits_{v\sim u}\frac{1}{w_{uv}}}-\kappa_{e}).$
@@
Given that $w_{e}(t)$ tends to infinity, there exists a sufficiently large time $t_{0}>0$ where $w_{e}^{\prime}(t_{0})>0$. Assume $w_{e}(t)$ does not have a monotonic increase on the interval $t_{0},\infty)$, then there exists $t_{1}>t_{0}$ such that $w_{e}^{\prime}(t_{1})<0$. By the continuity, there exists a point $a$ in $(t_{0},t_{1})$ such that $w_{e}^{\prime}(a)=0$. Given that $w_{e}(t)$ is always positive, we have that $\kappa_{e}(t_{0})<0$ and $\kappa_{e}(a)=0$.
@@
Notice that for any internal edge $uv$ and finite time $a$, we have $\kappa_{uv}(a)<0$. If $d_{u}^{\prime}=1$, there is only one leaf $e$, apparently $\kappa_{e}^{\prime}(a)<0$. Let $d_{u}^{\prime}>1$. W.l.o.g., we assume that $e$ attains the maximum initial weight among all leaves sharing $u$. Then for any leaf edge $g\sim e$, $\kappa_{g}(a)\leq 0$ according to Lemma 5. Since $D_{u}(t)$ is always positive, we conclude that $\kappa_{e}^{\prime}(a)<0$ according to the formula (7).
@@
Hence, for every leaf $e$, if its unnormalized weight grows to infinity, there exists some time $t_{0}$ such that $w_{e}^{\prime}(t)\geq 0$ for all $t>t_{0}$. The result follows.
@@ 4. Asymptotic behavior of curvature
Let the metric $w$ of T evolve under the Ricci flow (1). Consider $e$ as a leaf edge connected to vertex $u$ where $d_{u}\geq 3$. Consequently, the following statements are valid.
(i)  $d_{u}^{\prime}=d_{u}-1$ if and only if $\lim\limits_{t\to\infty}\kappa_{e}(t)=\frac{1}{d_{u}-1}$;
@@
WLOG, let $w_{e}(0)\geq w_{g}(0)$ for any leaf $g$ incident to $u$. When $d_{u}^{\prime}=d_{u}-1$, according to Lemma 5, there exists $t_{0}>0$ such that, $\kappa_{e}(t)\geq 1-\frac{d_{u}-2}{d_{u}-1}=\frac{1}{d_{u}-1}$ for $t\geq t_{0}$, which gives
@@
Conversely, assume that $\lim\limits_{t\to\infty}\kappa_{e}(t)=\frac{1}{d_{u}-1},$ it is easy to see that $\lim\limits_{t\to\infty}w_{e}(t)=0,$ therefore,
$\lim\limits_{t\to\infty}\frac{w_{e}}{w_{f}}(t)=0,$
for every internal edge $f\sim u$. Thus, $d^{\prime}_{u}=d_{u}-1.$ We proved the result (i).
@@
On the other hand, suppose that $d_{u}^{\prime}=d_{u}-2$. By Lemma 5, we have $\kappa_{e}(t)>0$ for large time $t$, which implies that $w_{e}(t)$ is decreasing. Consequently, for every internal edge $f\sim e$, the ratio $\frac{w_{e}}{w_{f}}$ is decreasing since the denominator is non-decreasing. It follows that
@@
which implies that $\frac{2-d_{u}}{w_{f}D_{u}}$ does not vanish along any time sequence, and hence neither does $\kappa_{f}$. On the other hand, since $\kappa_{e}$ remains positive for sufficiently large $t$, this contradicts the fact that $(\kappa_{f}-\kappa_{e})$ tends to zero.
Therefore, $\lim\limits_{t\to\infty}\frac{w_{e}}{w_{f}}=0$ for every internal edge, it follows that
@@
Let the metric $w$ of tree T evolve under the Ricci flow (1). Let $e=uv$ be a leaf with $d_{u}\geq 2$. Then the unnormalized weight $w_{e}(t)$ is bounded if and only if $d^{\prime}_{u}=d_{u}-2$ or $d^{\prime}_{u}=d_{u}-1$.
@@
By Proposition 3, the $\lim_{t\to\infty}w_{e}(t)$ exists. Suppose that $\lim_{t\to\infty}w_{e}(t)<\infty$. By Barbalat’s Lemma, the derivative
@@
must then converge to zero. Thus, if $\kappa_{e}(t)$ does not approach zero along at least one time sequence, then $w_{e}(t)\to 0$ along this sequence. Since the limit of $w_{e}(t)$ exists, we conclude that in fact
@@
Then, for each internal edge $f\sim e$, the ratio $\frac{w_{e}}{w_{f}}$ tends to zero as $t\to\infty$. Assume $d^{\prime}_{u}\leq d_{u}-3$, then we have
@@
Thus, $d^{\prime}_{u}$ is at least $d_{u}-2$ and $\kappa_{e}(t)$ approaches either $\frac{1}{d_{u}-1}$ or zero.
@@
Thus, the derivative $w^{\prime}_{e}(t)$ is negative, the unnormalized weight $w_{e}(t)$ decreases. Then $\lim\limits_{t\to\infty}w_{e}(t)$ is finite.
@@
(8) $\kappa_{e}\leq 1-\frac{d_{u}-2}{d_{u}^{\prime}+\frac{(d^{\prime\prime}_{u}-2)d^{\prime\prime}_{u}}{(d^{\prime\prime}_{u}-2+d_{u})}}.$
@@
According to the proof of Proposition 5, if $d^{\prime}_{u}\geq d_{u}-2$, as well as $d^{\prime}_{v}\geq d_{v}-2$, then there exist leaves attached to $u$ and $v$ such that the ratio of their unnormalized weights to the unnormalized weight of the edge $f$ approaches zero. Consequently, $\kappa_{f}$ tends to zero as $t$ approaches infinity as both $\kappa_{u\to f}\to 0$ and $\kappa_{v\to f}\to 0$ hold. ∎
@@ 5. The limit metric with constant curvature zero
Proof of result (3) Let $L$ be the set of leaf edges incident to $v_{1}$ or $v_{\ell}$. For finite $t$,
$\sum\limits_{\text{leaf~}e\sim v_{1}}\kappa_{e}(t)$ $=d_{v_{1}}-1-\frac{d_{v_{1}}-2}{D_{v_{1}}}\sum\limits_{\text{leaf~}e\sim v_{1}}\frac{1}{\tilde{w}_{e}}$
@@
$\sum\limits_{e\in L}\kappa_{e}(t)\downarrow 2,\quad t\to\infty,$
where $\downarrow 2$ means decrease to value $2$.
@@
The initial metric can influence the limiting weight value for a caterpillar tree. Consider a caterpillar tree where each spine vertex has a degree of $2$. Thus, leaf edges only incident to the endpoints of the main path.
@@
Repeating this argument for at least three different branches from $v_{0}$ shows that there are at least three such internal vertex $v$ in $T$ that are adjacent to $d_{v}-1$ leaves. ∎